Use the following information to answer these questions.

On January 1, our company purchased a truck for $85,000. The estimated useful life of the truck is 4 years. The residual value at the end of 4 years is estimated to be $5,000.

What is the depreciation expense for the second year of use if we use the double-declining balance method?

                           [ Select ]                       ["$40,000", "$22,500", "$21,250", "$42,500"]      

What is the balance in accumulated depreciation at the end of the second year of use if we use the double-declining balance method?

                           [ Select ]                       ["$63,750", "$62,500", "$42,500", "$45,000"]      

What is the book value at the end of the second year of use if we use the double-declining balance method?

                           [ Select ]                       ["$42,500", "$45,000", "$22,500", "$21,250"]      
Answer #1

Answer- The depreciation expense for the second year of use if we use the double-declining balance method = $21250.

The balance in accumulated depreciation at the end of the second year of use if we use the double-declining balance method = $42500+$21250

= $63750.

The book value at the end of the second year of use if we use the double-declining balance method = $21250.

Explanation- Double Declining balance depreciation is calculated using the following formula-

Depreciation = Depreciation Rate * Book Value of Asset

Depreciation rate is given by the following formula:

Depreciation Rate = Accelerator *Straight Line Rate

Straight-line Depreciation Rate = 1/4 = 0.25 = 25%
Declining Balance Rate = 2*25% = 50%

Depreciation for Year 1 = $ 85000*50% = $42500

Book value at the end of Year 1= $85000-$42500 =$42500

Depreciation for Year 2 = $ 42500*50% = $21250

Book value at the end of Year 2= $42500-$21250 =$21250